OpenStudy (moonlitfate):

Determine dy/dx for x*sin(y)=3

OpenStudy (moonlitfate):

The answer that I got for this was dy/dx = -cos y ; but, I'm not really sure how to do this.

OpenStudy (anonymous):

take the derivative of both sides, thinking that \(y=f(x)\) so you need both the product and the chain rule for the left the derivative of the right is 0

OpenStudy (anonymous):

\[x\sin(y)=3\] \[\sin(y)+x\cos(y)y'=0\]

OpenStudy (anonymous):

solve the equation for \(y'\) via \[x\cos(y)y'=-\sin(y)\] \[y'=-\frac{\sin(y)}{x\cos(y)}\]

OpenStudy (anonymous):

like being asked for the derivative of \[x\sin(f(x))\] need both product and chain rule

OpenStudy (moonlitfate):

@satellite73-- What if I had something like:2x^2+x*y+3y^2=6 ?

OpenStudy (anonymous):

then you need the product and chain rule again like being asked for the derivative of \[2x^2+xf(x)+3f^2(x)\]

OpenStudy (moonlitfate):

Hmm. So, wherever I have a y in the equation, I can replace it with f(x)...

OpenStudy (anonymous):

the derivative of \(2x^2\) is \(4x\) the derivative of \(xf(x)\) is \(f(x)+xf'(x)\) by the product rule and the derivative of \[3f^2(x)\] is \(6f(x)f'(x)\) again by chain rule

OpenStudy (anonymous):

yeah that is my explanation not really what you want to do you are thinking that \(y\) is a function of \(x\) even though you have not written it explicitly as one that is why it is called "implicit" differentiation

OpenStudy (anonymous):

\[2x^2+xy+3y^2=6 \] \[4x+y+xy'+6yy'=0\] is a lot easier to write, than it is to replace \(y\) by \(f(x)\) and \(y'\) by \(f'(x)\)

OpenStudy (moonlitfate):

That's what I had before with the y'

OpenStudy (anonymous):

good

OpenStudy (moonlitfate):

Would the answer be: -5x/6y? :O

OpenStudy (anonymous):

oh no requires more algebra than that

OpenStudy (moonlitfate):

Aw, I got excited for a minute.

OpenStudy (anonymous):

\[4x+y+xy'+6yy'=0\] \[xy'+6yy'=-4x-y\] \[(x+6y)y'=-4x-y\] \[y'=\frac{-4x-y}{x+6y}\]

OpenStudy (anonymous):

it is just algebra though not calculus at that step

OpenStudy (moonlitfate):

Ohh, so the derivative of whatever variable I'm taking stays to the left?

OpenStudy (anonymous):

you have to solve for \(y'\) using algebra, no matter where it is

OpenStudy (anonymous):

\[x^2+y^2=xy\] \[2x+2yy'=y+xy'\] you still have to solve for \(y'\)